Directions:Solve this system of equations by graphing.
{-4x+2y=8
{6x-3y=3

Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39618)

You can't.
Already responded.
Compare 4/2=2 with 6/3=2.

System, yes;
Solve, no.

Answer by MathTherapy(10552)

Directions:Solve this system of equations by graphing.
{-4x+2y=8
{6x-3y=3

NO SOLUTION, since the graphs of the equations are parallel lines. You should graph them, and you'll see that
the lines will NEVER intersect since they're parallel. Additionally, the graphs have the same slope: 2.